OpenStudy (anonymous):

f(x) =4/x + 5 & g(x) =x/x + 5 what is f+g?

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x)=4/(x+5)+x/(x+5)=(4+x)/(x+5)

OpenStudy (anonymous):

so all i do it add the two?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you very much.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but how did you get (4+x)/(x+5) wouldnt it be just 4x/x+5 ?